As in the previous question, suppose you have made 1L worth of a mixture of methanol and water, by mixing 3 parts of methanol and 2 parts of water by mass. Assume for simplicity that the volume of the mixture is exactly equal to the sum of the volumes of the ingredients (which is not true, by the way). What is the mass concentration of oxygen in the mixture? (Note that the vast majority of oxygen is bound to other atoms.) Give answer in g/L, 2 significant digits, do not write down the units.
Answer #1

From the question

1L solution has 1 kg mass

And

We know that

1kg = 1000g

Let mass of methanol = 3x and mass of water = 2x

Because

Mass of methanol + mass of water = 1000g

So

3x + 2x = 1000g

5x = 1000g

x = 1000g/5

x = 200g

mass of methanol = 3×200g = 600g

mass of water = 2×200g = 400